the magnifying glass is made of lenses of power +20D and -4D. if distance of distinct vision is 25cm. calculate the size of an object 2cm high seen through the magnifying glass
1
Expert's answer
2020-04-29T09:48:31-0400

As per the given question,

The power of the lenses are "(P_1)=+20D" and "(P_2)=-4D"

distance of distinct vision (v) =25cm

The size of the object "(h_o)=2cm"

Let the focal length of the glasses are "f_1" and "f_2"

We know that,

"f_1=\dfrac{1}{P_1}=\dfrac{100}{20}=5cm"

Similarly,

"f_2=\dfrac{1}{P_2}=\dfrac{-100}{4}=-25cm"

Now, combined focal length "\dfrac{1}{f}=\dfrac{1}{f_1}+\dfrac{1}{f_2}=\dfrac{1}{5}-\dfrac{1}{25}=\dfrac{4}{25}"

"f=\dfrac{25}{4}"

Now, let the distance of object is u

"\Rightarrow \dfrac{1}{v}-\dfrac{1}{u}=\dfrac{1}{f}"

"\Rightarrow \dfrac{1}{u}=\dfrac{1}{25}-\dfrac{4}{25}=\dfrac{3}{25}"


So, "u=\dfrac{25}{3}cm"

we know that,

"\dfrac{h_i}{h_o}=\dfrac{-v}{u}"

"h_i=\dfrac{25\times 3\times 2}{25}=-6cm"